OpenStudy (anonymous):

Find the value of 3/ sqrt 3? 3 3 sqrt 3 sqrt 3 sqrt 3 / 3

OpenStudy (mayankdevnani):

\[\huge \frac{3}{\sqrt{3}}\] first, multiply sqrt{3} on both numerator and denominator. like, \[\huge \frac{3 \times \sqrt{3}}{\sqrt{3} \times \sqrt{3}}\] and then simply multiply it

OpenStudy (mayankdevnani):

what you get when you solve it? @Katessxo

OpenStudy (anonymous):

its root3

OpenStudy (rsadhvika):

yes

OpenStudy (mayankdevnani):

@rathanreddy plzz don't give direct answers. let @Katessxo think

OpenStudy (anonymous):

3 can be written as root3*root3 the root3 in the denominator and numerator are cancelled

OpenStudy (anonymous):

so, root3 is answer

OpenStudy (mayankdevnani):

use EQUATION EDITOR.... @rathanreddy

OpenStudy (mayankdevnani):

so @Katessxo can you solve it???

OpenStudy (anonymous):

Oh okay i understand now. Thank you both for your help!

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (rsadhvika):

an easy way what rathan was saying is :- \(\large \frac{3}{\sqrt3}\) \(\large \frac{\sqrt{3} \times \sqrt{3}}{\sqrt3}\) \(\large \frac{\cancel{\sqrt{3}} \times \sqrt{3}}{\cancel{\sqrt3}}\) \(\large \sqrt{3}\)

OpenStudy (anonymous):

Thats what i figured but it still was a bit hard to understand. Thank you @rsadhvika

OpenStudy (rsadhvika):

np :D

OpenStudy (anonymous):